OpenStudy (anonymous):

help? A plumber finished three jobs on Tuesday. The first two only cost the owner the $45 trip fee because they took very little time to complete. For the third job, the plumber charged the trip fee plus 6 times his hourly rate. If the plumber received a total of $303 for the day, what is the hourly rate?

OpenStudy (swissgirl):

ok so we know that he recieved $45 for the first 2 jobs so that is 45+45=$90 Now the third job he charged $45 plus 6 times his hourly rate which is 45+6x Now in total he received 90+45+6x=303 Your x would be the hourly rate. So you gotta solve for x

OpenStudy (anonymous):

okay.

OpenStudy (swissgirl):

135+6x=303 6x=168 x=28 So his hourly rate is $28

OpenStudy (anonymous):

Yeah, I just did it.

OpenStudy (swissgirl):

YAYYYYYY:)

OpenStudy (anonymous):

YEEEAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAA lol.